Why can't the angle between force and displacement be 270 degrees in case of motion in a plane
1
Expert's answer
2020-11-12T09:47:04-0500

"\text{If the angle } \alpha \text{ between vector and axis, then the projection of force}"

"\text { along axis is found as }\vec{F{pr}}=\vec{F}*cos{\alpha}"

"\vec{F{pr}}=\vec{F}*cos{\alpha}=\vec{F}*cos{270^0}=0"

"|\vec{F{pr}}|=0"

"\text{Therefore, the body does not move at this angle.}"

Answer: The body does not move at an angle of 270 to the applied force.